We have a closed system with an aluminum cup, water and a cube of ice. The glass has a mass of 100 grams. The temperature of the glass and water is 20 degrees C. We also have 200 grams of water in total. The cube of ice weighs 5 grams and is at 0 degrees C. Find the equilibrium temperature of the system.
Answer #1

mass of water = was ; abot 927-2 (20-T) a, = 13544-927-2T louo 2 = 1670十2093T ol en | 670tga, 95T = 18544-927-27 (D.qs+ 90.1216874 94313 T- 17.8。